OpenStudy (anonymous):

SHOWYOURWORK! (2x^2)(-8x^3) and solve the system of equations 2x-y=-10 3x+4y=1

sam (.sam.):

There's 2 questions there?

sam (.sam.):

(2x^(2))(-8x^(3)) Multiply 2x^(2) by -8x^(3) to get -16x^(5). -16x^(5)

OpenStudy (anonymous):

and then solve the two last eqautions

sam (.sam.):

2x-y=-10..................(1) 3x+4y=1 ...............(2) From 2x-y=-10, 2x=y-10 x=(y-10)/2 ------------------------ From (2), 3x+4y=1 3(y-10)/2+4y=1 Multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of 2. (3(y-10))/(2)+4y*(2)/(2)=1 Multiply 4y by 2 to get 8y. (3(y-10))/(2)+(8y)/(2)=1 The numerators of expressions that have equal denominators can be combined. In this case, (3(y-10))/(2) and ((8y))/(2) have the same denominator of 2, so the numerators can be combined. (3(y-10)+(8y))/(2)=1 Combine all similar expressions in the polynomial. (11y-30)/(2)=1 Multiply each term in the equation by 2. 11y-30=2 move all terms not containing y to the righthand side of the equation. 11y=32 Divide each term in the equation by 11. y=32/11 ------------------------------------------------- x=(y-10)/2 x=((32/11)-10)/2 y=-(39)/(11)

sam (.sam.):

x=-(39)/(11) , not y